Circular No. 7572

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2001 B2 S. H. Pravdo, Jet Propulsion Laboratory, reports the discovery by

NEAT of an apparent comet. Available astrometry:

2001 UT α_{2000} δ_{2000} m_{1} Observer
Jan. 24.60450 12^{h}04^{m}56.^{s}15 −22^{o}05′45′′.6 15.5 NEAT
24.61471 12 04 55.53 −22 05 43.4 15.4  _{”}
24.62490 12 04 55.07 −22 05 40.7 15.1  _{”}
25.29734 12 04 22.67 −22 03 00.9  Lopez
25.30243 12 04 22.37 −22 02 58.9  _{”}
25.31493 12 04 21.93 −22 02 59.2 14.7 Durig
25.31649 12 04 21.73 −22 02 55.1  Lopez
25.33038 12 04 21.04 −22 02 52.0  _{”}
25.34427 12 04 20.36 −22 02 48.3  _{”}
25.35451 12 04 19.82 −22 02 51.2 14.6 Durig
25.37222 12 04 18.99 −22 02 45.5 14.2  _{”}

NEAT (Haleakala, HI). Observers E. F. Helin, S. Pravdo, K. Lawrence, P.

Kervin, R. Maeda, and M. Skinner. 1.2-m reflector. C. E. Lopez (San Juan, Argentina). 0.5-m f/7.5 double astrograph. D. T. Durig (Sewanee, TN). 0.3-m f/5.75 reflector. Diffuse tail ∼ 25′′ long

toward the south-southeast.

SUNGRAZING COMETS Further to IAUC 7567, D. Hammer reports his measurements for sev-

eral additional Kreutz sungrazing comets from SOHO C2 coronagraph data. Comets C/2000 H5 and C/2000 X8 were found by D. Biesecker, and the remaining comets were found at the SOHO website by M. Oates. The reduced observations and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2000 UT  α_{2000}  δ_{2000}  MPEC
C/2000 H3 Apr. 17.254  1^{h}49.^{m}4  + 9^{o}40′  2001-B21
C/2000 H4  30.504  2 37.7  +13 50  2001-B21
C/2000 H5  30.504  2 37.8  +13 49  2001-B21
C/2000 J7 May 4.338  2 51.6  +14 50  2001-B12
C/2000 L6 June 11.229  5 16.7  +21 22  2001-B12
C/2000 N3 July 4.785  6 50.4  +21 47  2001-B12
C/2000 X8 Dec. 12.743  17 26.0  −25 01  2001-B12
2001 January 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green